OpenStudy (anonymous):

The Dead Sea has an elevation of one thousand three hundred ten feet below sea level. Mount Everest has an elevation of twenty-nine thousand twenty-eight feet above sea level. How many feet higher is Mount Everest than the Dead Sea?

OpenStudy (anonymous):

1310+29028=30338 ft

OpenStudy (plainntall):

It looks like you are learning negative numbers. One way to thing of elevation is by taking the difference of the highest value and the lowest value. The value for Everest is 29,028 ft, sea level is 0 and the Dead Sea is -1310 ft When we want to find the distance between the top of Everest and the bottom of the Dead Sea we subtract their values. Distance = Everest - Dead Sea = 29,028 - (-1310) What happens when we multiply a negative number by a negative number? It becomes a positive so. =29.028 +1310 = 30,338 as Yoloswagoo said. The tricky part is to know the formula for distance D2-D1 and that multiplying a negative number by -1 changes it to a positive number.